OpenStudy (lukecrayonz):

A rectangular plot of farmland will be bounded on one side by a river, and on the other three sides by a single strand electric fence. With 1900m of wire at your disposal, what is the largest area you can enclose, and what are its dimensions?

OpenStudy (anonymous):

let a be the length and b be the breadth so. area=A=ab ........(1) 2a+b=1900 i.e b=1900-2a .........(2) sustitute eq one in two A=a(1900-2a) A=1900a-2a^2 Now find dA/da dA/da=1900-4a For maximum area put dA/da=0 hence 0=1900-4a a=475m so b=1900-475=1425m You will have maximum area in dimentions 475x1475